# Two Shades of Quark Color: Parallel Canons across the Cold War Divide

### v1

@@ Abstract
The introduction of the color quantum number is conventionally narrated as a linear progression from the quark-model statistics paradox to quantum chromodynamics (QCD). This paper challenges that teleology by arguing that “color” emerged as two conceptually distinct constructs during the Cold War. The first, originating with Han and Nambu and culminating in QCD, conceived of color as a local gauge charge, the source of a fundamental force mediated by gluons. The second, developed at the Joint Institute for Nuclear Research (JINR) in Dubna, treated color as a hidden, three-valued label—a statistical and structural property within a composite, S-matrix-inflected hadron model. We trace these parallel narratives, linking the Dubna approach to a holist epistemology that prioritizes observable amplitudes and global constraints, and the QCD approach to a reductionist program grounded in micro-dynamics. A case study of Fermilab’s E–36 experimental chain (1970–78) shows how an observables-first design—tuned to S-matrix and Regge constraints on forward elastic scattering—performed robustly within its natural domain but was ultimately discontinued amid declining theoretical interest and involvement. The subsequent hegemony of QCD retroactively projected its gauge-theoretic conception of color onto history, erasing this epistemic diversity. We conclude that the marginalization of Dubna’s structural color was not merely a political outcome of the Cold War but a result of deep ontological and philosophical divergences, advocating for a domain-sensitive pluralism in the historiography of particle physics.
@@ 1 Introduction
The introduction of the color quantum number is often taken—retrospectively—as the hinge between phenomenological schemes for organizing hadrons and a dynamical field theory of their constituents, quantum chromodynamics (QCD) [8]. The canonical storyline runs from the quark model’s statistics paradoxes [1, 2] through the postulation of an additional internal degree of freedom [3, 4, 5] to the maturation of a non- Abelian gauge account in which color is a local charge [6, 7, 8, 9, 10]. This paper resists that teleology. We argue that “color” emerged as two conceptually distinct constructions: a hidden, three-valued label introduced within composite-quark phenomenology (the Dubna line), and a local gauge charge in Yang–Mills theory (the Han–Nambu thread and the later QCD synthesis). These constructions grew in partially overlapping but differently normed research cultures, partially incommensurable and yielding parallel —sometimes rival—epistemic orientations rather than a single, linear path to QCD.
@@ 2.1 Dubna’s “color as label” (1965)
In early 1965, Dubna authors proposed an additional internal quantum number for quarks—what would soon be called color—precisely to reconcile the SU(6) quark spectroscopy with Fermi–Dirac/Pauli constraints in the composite quark model frame [4, 5, 19, 11]. In this narrative, color first functions as a hidden index that antisymmetrizes the otherwise symmetric spin–flavor wavefunctions and enforces selection rules at the level of residues in amplitudes, while hadrons observed in nature are taken to be color neutral (singlets) [11, 20]. The JINR work of 1965 develops a dynamic composite-hadron model in which quarks are degrees of freedom yet permanently confined, and only color-singlet hadrons appear as asymptotic states (the free, experimentally detectable particles (like hadrons) long before or after an interaction, which form the basis of the S-matrix formalism) [20]. Color here labels unitary-equivalent internal states that do bookkeeping in the residue algebra, the mathematical structure governing the strengths (residues) of bound-state poles in scattering amplitudes, encoding selection rules and symmetry constraints. No explicit gluon dynamics is posited at this stage.
Historically, Struminsky’s 1965 Dubna work introduced an additional internal quantum number for quarks to reconcile SU(6) with Fermi–Dirac statistics; it did not posit an SU(3) local gauge theory or an octet of gluons [5]. The first move to put the new degree of freedom on an SU(3) gauge footing—explicitly with eight vector fields as interaction carriers— was by Han and Nambu in 1965 [3]; see also Matveev–Tavkhelidze’s retrospective noting Nambu’s eight vectors as the gluon prototype [11]. The parafermi statistics route, introduced by Greenberg[21], provided a global, hidden degree of freedom. While it could later be connected to a gauge theory—for instance, via the Green ansatz leading to an SO(3) model with three gauge bosons, which is phenomenologically distinct from QCD—its original formulation was not that of a local gauge theory. Despite Struminsky being first to use an SU(3) symmetry in color studies, in his formulation it served as a global SU(3)–type label (three unitary-equivalent quark states) rather than a local gauge symmetry [5, 11]. The parafermi route was later shown to be equivalent to the SU(3) color model in its classification of the allowed color-singlet hadronic states, not as a dynamically equivalent SO(3) gauge theory [22, 23].
The Dubna program subsequently developed the ”quasi-free quark model” where color neutrality ($\epsilon^{\alpha\beta\gamma}$ for baryons, $\delta_{\alpha}^{\beta}$ for mesons) defined asymptotic states while quarks remained permanently confined through dynamical mechanisms external to the color concept itself [11, pp. 308–309].
@@
These parallel developments reveal the period’s conceptual fluidity: identical phenomenological constraints (statistics, confinement) admitted multiple physically distinct solutions, with the choice between them reflecting deeper commitments to either composite models or emergent gauge theories.
@@
Secondly, sum rules function as normative constraints. Current algebra with dispersive integrals (Adler; Adler–Weisberger) ties symmetry commitments to measurable rates [35]. Philosophically, these are not discoveries of hidden objects but obligations our amplitude calculus must satisfy if the symmetry talk is to earn its keep. The Shifman–Vainshtein–Zakharov SVZ (IHEP, Protvino) sum rules played a role of an accountability ledger. The SVZ program (OPE + dispersion) relates “micro” coefficients (condensates in operator products) to “macro” spectral weights of hadrons [36]. Color remains absent from the S-matrix spectrum; what is present is a disciplined matching between two descriptions—one written in operator coefficients, the other in colorless spectral functions.
@@ 2.5 What’s at Stake Conceptually
The essential distinction between the two color concepts can be captured aphoristically: Dubna “color” = symmetry of amplitude/residue; QCD color = source of the force. In the 1965 Dubna scheme, quarks are treated as confined constituents (entity-level commitment), while “color” functions as a global SU(3)-type index that stabilizes the residue algebra (Pauli/SU(6) consistency and selection rules). Thus color is real as a structural constraint, not as a causally efficacious charge. The subsequent QCD consolidation refunctionalizes the same SU(3) into a local gauge symmetry: color becomes a dynamical charge (with gluon carriers), i.e., real as a source term in the field equations. In short: quarks-as-entities with color-as-constraint (Dubna), versus quarks-as-entities with color-as-cause (QCD).
In the Dubna framework, color governed the combinatorial structure of hadronic wavefunctions and selection rules—a property of composite systems rather than fundamental fields. As Matveev emphasizes, their approach treated ”colored quarks as real physical objects defining the hadron structure” within a dynamical model where confinement was presupposed rather than explained [11, pp. 307–308].
@@
This distinction maps onto a broader epistemic divide: compositeness models treat properties as emergent from symmetry and structure, while gauge theories derive them from localized interaction mechanisms encoded in field dynamics. Struminsky models baryons as composites of three quarks and then adds an additional quantum number so that the residue algebra respects Pauli/SU(6) without ever populating the asymptotic spectrum with colored states [5, 4]. In short, quarks are the constituent variables of a composite description, while color is a global SU(3)-type index that makes the whole amplitude narrative cohere—precisely a holist use of structure, not yet endowed with genuine dynamics. While not derived from S-matrix theory, Struminsky’s approach reflects its holist spirit: it privileges observable, colorless states (only colorless hadrons appear as $S$-matrix states); treats color as a structural constraint rather than a gauge-mediated force; stabilizes the amplitude bookkeeping rather than a local cause; emphasizes internal consistency over microscopic causation—placing it epistemologically closer to compositional holism than to dynamical field theory.
@@ 3 Holism and reductionism in particle physics
This exemplifies what philosophers call entity realism—the position that we can be realist about entities we can causally manipulate, even when their theoretical description remains incomplete [44, 45]. Holism treats structures, constraints, and organizing relations as explanatorily primary. The ancient template is already clear: Plato’s Timaeus makes order-by-form the driver of explanation (wholes first, parts second), while the Stoics’ continuous cosmos—held together by logos/pneuma—models nature as a field-like unity rather than discrete bits [46, 47]. Aristotle’s hylomorphism (form+matter) likewise takes organized wholes to enjoy a kind of explanatory priority over mere aggregates [48]. In modern physics, this idiom reappears whenever global principles and symmetry/analytic constraints carry more weight than microscopic bookkeeping.
A mid-20th-century expression is the S-matrix/bootstrap program: start from analyticity, unitarity, and crossing, and demand a self-consistent network of colorless amplitudes—no micro ontology required [15, 14]. Philosophically, that’s close to structural realism: what survives—and does the explaining—are relations, symmetries, and constraints, not independently grasped “things” [49, 50, 51]. In this key, “color” can function as a constitutive symmetry of the residue algebra (selection rules, singletness) without appearing as an asymptotic quantum number—precisely the Dubna narrative.
@@
This structuralist stance treats physical theories not as descriptions of fundamental entities but as networks of constraints that generate observable phenomena—a view that finds natural expression in both the bootstrap program and the Dubna approach to color.
@@
The USSR and pre-USSR philosophical tradition includes holist frameworks that, while not direct influences, conceptually resonate with structural approaches in physics. These frameworks emphasize explanation via global constraints, symmetry, and system-level organization—features aligned with the Dubna school’s early non-dynamical conception of quark color. For example, Vladimir Vernadsky treated the biosphere as a planetary-scale system whose constraints govern local dynamics. His notion of the noosphere—human thought as a geophysical force—anticipates modern views of scale-dependent structure. Vernadsky’s framework aligns with effective field theory and renormalization group approaches in its holist orientation: macro-level patterns exhibit autonomy from micro-details, and explanation centers on the global constraint architecture that renders large-scale behavior intelligible [56].
These philosophical traditions should not be read as methodological prescriptions that directly determined Dubna’s scientific approach. An example is the Bogolyubov–Shirkov renormalization group [69], grounded in axiomatic quantum field theory, which reflects a reductionist orientation: it formalizes how local interaction structures constrain behavior across energy scales, aiming to preserve consistency from micro-level dynamics upward—yet its emphasis on symmetry and analyticity also lends it a structural, quasi-holist dimension.
Rather, holist frameworks represent the broader intellectual ecosystem in which Soviet theoretical physics developed during the Cold war. Scientific communities, while governed by empirical and mathematical constraints, nevertheless operate within cultural contexts that can shape methodological preferences and theoretical sensibilities. The presence of these holist traditions in USSR thought may have helped cultivate an intellectual environment where constraint-based, structural explanations felt particularly natural—not as a matter of philosophical determinism, but as part of the conceptual resources available to physicists thinking through fundamental problems. This cultural context helps explain why the Dubna approach to color, with its emphasis on structural constraints over dynamical charges, emerged as a coherent alternative to the gauge-theoretic paradigm developing elsewhere.
@@ 4.1 Problem choice and ontology
The E–36 experiment and its chain were designed to interrogate the high–energy, small–$|t|$ corner of elastic $pp$ scattering where the native theoretical grammar is the Regge/S-matrix one, not perturbative QCD. Already in the early 1970s—and still today for soft elastic diffraction—perturbative QCD offers no general, first-principles description of forward amplitudes, whereas the Regge framework provides a compact language of analytic constraints, pole/trajectory structure, and factorization tailored to this regime. [^1] On the domain-limited tractability of perturbative QCD for small-angle scattering and the Regge/S-matrix idiom used by the E–36 collaboration, see the historical analysis in Phys. in Perspective 18 (2016) 357–378, esp. the section “Experiment Ahead of QCD.” In Regge theory unknown amplitudes are parametrized by Regge trajectories (Reggeons) and their residues; the vacuum exchange that organizes high–energy diffraction is the pomeron.[57, 16, 58, 59, 60, 61] It was precisely this observables-first ontology—analytic $S$ with poles/cuts, crossing, unitarity—that E–36 was built to test.
@@ 4.2 The E36 experimental design as a realization of the S-matrix program
The technological development itself was theory-laden: K.D. Tolstov’s cryogenic target design at JINR emerged from research priorities set by Pomeranchuk’s work on high-energy scattering theory. The target’s optimization for studying elementary proton interactions rather than searching for new particles reflected the S-matrix tradition’s focus on fundamental scattering processes over particle discovery.
@@
While later QCD reinterprets the Pomeron as a color-singlet gluon ladder, the E–36 chain neither requires nor discriminates that ontology; its evidential force lies in validating amplitude-level constraints that the Dubna program treated as primary.
@@ 4.3 Results as S-matrix constraints
E–36’s initial Physical Review Letters reported (a) a rise of $B(s)$ with energy across 88400400 GeV and (b) a transition to positive $\rho$ in the few-hundred GeV range. [^5] PRL slope and $\rho$ papers: Phys. Rev. Lett. 31 (1973) 1088–1091; 31 (1973) 1367–1370. The $\rho>0$ finding around $\sim\!300$ GeV catalyzed a re-reading of the Pomeranchuk theorem among practitioners and clarified how forward analyticity constraints should be applied in practice; [^6] For discussion of the $\rho$ sign change and its interpretive consequences, see [63, 16]. together with the measured shrinkage, these data were absorbed naturally in the Regge phenomenology that motivated the experiment. [^7] On the centrality of Regge/Pomeron language in E–36 analysis and the enduring absence of a first-principles QCD derivation of Reggeons, see [16, 60, 61].
Crucially, the collaboration could—and did—interpret its forward data entirely within the S-matrix/Regge idiom, without invoking quarks or gluons.888“The participants in E–36 and its subsequent chain of experiments did not have to refer to QCD or the quark model in their papers in order to interpret their results.” [62, 63, 64, 65] This was not a deficiency but a principled alignment of domain and ontology: when the task is to constrain $f(s,t)$ near $t=0$, the right language is the one that treats $f$ as primary.
@@
E–36 inhabited the soft, small–$|t|$ domain where QCD’s perturbative “crispness” is inapplicable and where QCD—then nascent—offered neither mechanism nor method for diffraction. In that domain, explanation comes from global structure: the pole content and analyticity of $S$, Regge shrinkage, and the high–energy behavior codified by the pomeron. [^9] On the historical timing (E–36 underway prior to asymptotic freedom) and on the acknowledged gap between soft Regge behavior and QCD derivations, see [65, 64, 9, 10] and [16, 60, 61]. Treating forward constraints as first-class physics—and building an apparatus around them—is non-reductionist in the precise sense relevant to the S-matrix tradition.

### v2

@@ Abstract
The introduction of the color quantum number is conventionally narrated as a linear progression from the quark-model statistics paradox to quantum chromodynamics (QCD). This paper challenges that teleology by arguing that “color” emerged as two conceptually distinct constructs during the Cold War. The first, originating with Han and Nambu and culminating in QCD, conceived of color as a local gauge charge, the source of a fundamental force mediated by gluons. The second, developed at the Joint Institute for Nuclear Research (JINR) in Dubna, treated color as a hidden, three-valued label—a statistical and structural property within a composite, S-matrix-inflected hadron model. We trace these parallel narratives, linking the Dubna approach to a holist epistemology that prioritizes observable amplitudes and global constraints, and the QCD approach to a reductionist program grounded in micro-dynamics. A case study of Fermilab’s E–36 experimental chain (1970–78) shows how an observables-first design—tuned to S-matrix and Regge restrictions on forward elastic scattering—performed robustly within its natural domain but was ultimately discontinued amid declining theoretical interest and involvement. The subsequent hegemony of QCD retroactively projected its gauge-theoretic conception of color onto history, erasing this epistemic diversity. We conclude that the marginalization of Dubna’s structural color was not merely a political outcome of the Cold War but a result of deep ontological and philosophical divergences, advocating for a domain-sensitive pluralism in the historiography of particle physics.
@@ 1 Introduction
The introduction of the color quantum number is often taken—retrospectively—as the hinge between phenomenological schemes for organizing hadrons and a dynamical field theory of their constituents, quantum chromodynamics (QCD) [9]. The canonical storyline runs from the quark model’s statistics paradoxes [1, 2] through the postulation of an additional internal degree of freedom [3, 4, 5] to the maturation of a non- Abelian gauge account in which color is a local charge [7, 8, 9, 10, 11]. This paper resists that teleology. We argue that “color” emerged as two conceptually distinct constructions: a hidden, three-valued label introduced within composite-hadron phenomenology (the Dubna line), and a local gauge charge in Yang–Mills theory (the Han–Nambu thread and the later QCD synthesis). These constructions grew in partially overlapping but differently normed research cultures, partially incommensurable and yielding parallel —sometimes rival—epistemic orientations rather than a single, linear path to QCD.
@@ 2.1 Dubna’s “color as label” (1965)
In early 1965, Dubna authors (see Endnote 1) proposed an additional internal quantum number for quarks—what would soon be called color—precisely to reconcile the SU(6) quark spectroscopy with Fermi–Dirac/Pauli constraints in the composite hadron model frame [4, 5, 20, 12]. In this narrative, color first functions as a hidden index that antisymmetrizes the otherwise symmetric spin–flavor wavefunctions and enforces selection rules at the level of residues in amplitudes, while hadrons observed in nature are taken to be color neutral (singlets) [12, 21]. The JINR work of 1965 develops a dynamic composite-hadron model in which quarks are degrees of freedom yet permanently confined, and only color-singlet hadrons appear as asymptotic states (the free, experimentally detectable particles (like hadrons) long before or after an interaction, which form the basis of the S-matrix formalism) [21]. Color here labels unitary-equivalent internal states that do bookkeeping in the residue algebra, the mathematical structure governing the strengths (residues) of bound-state poles in scattering amplitudes, encoding selection rules and symmetry requirements. No explicit gluon dynamics is posited at this stage.
Historically, Struminsky’s 1965 Dubna work (see Endnote 2) introduced an additional internal quantum number for quarks to reconcile SU(6) with Fermi–Dirac statistics; it did not posit an SU(3) local gauge theory or an octet of gluons [5]. The first move to put the new degree of freedom on an SU(3) gauge footing—explicitly with eight vector fields as interaction carriers— was by Han and Nambu in 1965 [3]; see also Matveev–Tavkhelidze’s retrospective noting Nambu’s eight vectors as the gluon prototype [12]. The parafermi statistics route, introduced by Greenberg[22], provided a global, hidden degree of freedom. While it could later be connected to a gauge theory—for instance, via the Green ansatz leading to an SO(3) model with three gauge bosons, which is phenomenologically distinct from QCD—its original formulation was not that of a local gauge theory. Despite Struminsky being first to use an SU(3) symmetry in color studies, in his formulation it served as a global SU(3)–type label (three unitary-equivalent quark states) rather than a local gauge symmetry [5, 12]. The parafermi route was later shown to be equivalent to the SU(3) color model in its classification of the allowed color-singlet hadronic states, not as a dynamically equivalent SO(3) gauge theory [24, 25].
The Dubna program subsequently developed the ”quasi-free quark model” where color neutrality ($\epsilon^{\alpha\beta\gamma}$ for baryons, $\delta_{\alpha}^{\beta}$ for mesons) defined asymptotic states while quarks remained permanently confined through dynamical mechanisms external to the color concept itself [12, pp. 308–309].
@@
These parallel developments reveal the period’s conceptual fluidity: identical phenomenological requirements (statistics, confinement) admitted multiple physically distinct solutions, with the choice between them reflecting deeper commitments to either composite models or emergent gauge theories.
@@
Secondly, sum rules function as normative criteria. Current algebra with dispersive integrals (Adler; Adler–Weisberger) ties symmetry commitments to measurable rates [37]. Philosophically, these are not discoveries of hidden objects but obligations our amplitude calculus must satisfy if the symmetry talk is to earn its keep. The Shifman–Vainshtein–Zakharov SVZ (ITEP, Moscow) sum rules played a role of an accountability ledger. The SVZ program (OPE + dispersion) relates “micro” coefficients (condensates in operator products) to “macro” spectral weights of hadrons [38]. Color remains absent from the S-matrix spectrum; what is present is a disciplined matching between two descriptions—one written in operator coefficients, the other in colorless spectral functions.
@@ 2.5 What’s at Stake Conceptually
The essential distinction between the two color concepts can be captured aphoristically: Dubna “color” = symmetry of amplitude/residue; QCD color = source of the force. In the 1965 Dubna scheme, quarks are treated as confined constituents (entity-level commitment), while “color” functions as a global SU(3)-type index that stabilizes the residue algebra (Pauli/SU(6) consistency and selection rules). Thus color is real as a structural condition, not as a causally efficacious charge. The subsequent QCD consolidation refunctionalizes the same SU(3) into a local gauge symmetry: color becomes a dynamical charge (with gluon carriers), i.e., real as a source term in the field equations. In short: quarks-as-entities with color-as-constraint (Dubna), versus quarks-as-entities with color-as-cause (QCD).
In the Dubna framework, color governed the combinatorial structure of hadronic wavefunctions and selection rules—a property of composite systems rather than fundamental fields. As Matveev emphasizes, their approach treated ”colored quarks as real physical objects defining the hadron structure” within a dynamical model where confinement was presupposed rather than explained [12, pp. 307–308].
@@
This distinction maps onto a broader epistemic divide: compositeness models treat properties as emergent from symmetry and structure, while gauge theories derive them from localized interaction mechanisms encoded in field dynamics. Struminsky models baryons as composites of three quarks and then adds an additional quantum number so that the residue algebra respects Pauli/SU(6) without ever populating the asymptotic spectrum with colored states [5, 4]. In short, quarks are the constituent variables of a composite description, while color is a global SU(3)-type index that makes the whole amplitude narrative cohere—precisely a holist use of structure, not yet endowed with genuine dynamics. While not derived from S-matrix theory, Struminsky’s approach reflects its holist spirit: it privileges observable, colorless states (only colorless hadrons appear as $S$-matrix states); treats color as a structural criterion rather than a gauge-mediated force; stabilizes the amplitude bookkeeping rather than a local cause; emphasizes internal consistency over microscopic causation—placing it epistemologically closer to compositional holism than to dynamical field theory.
@@ 3 Holism and reductionism in particle physics
This exemplifies what philosophers call entity realism—the position that we can be realist about entities we can causally manipulate, even when their theoretical description remains incomplete [46, 47]. Holism treats structures, requirements, and organizing relations as explanatorily primary. The ancient template is already clear: Plato’s Timaeus makes order-by-form the driver of explanation (wholes first, parts second), while the Stoics’ continuous cosmos—held together by logos/pneuma—models nature as a field-like unity rather than discrete bits [48, 49]. Aristotle’s hylomorphism (form+matter) likewise takes organized wholes to enjoy a kind of explanatory priority over mere aggregates [50]. In modern physics, this idiom reappears whenever global principles and symmetry/analytic conditions carry more weight than microscopic bookkeeping.
A mid-20th-century expression is the S-matrix/bootstrap program: start from analyticity, unitarity, and crossing, and demand a self-consistent network of colorless amplitudes—no micro ontology required [16, 15]. Philosophically, that’s close to structural realism: what survives—and does the explaining—are relations, symmetries, and conditions, not independently grasped “things” [51, 52, 53]. In this key, “color” can function as a constitutive symmetry of the residue algebra (selection rules, singletness) without appearing as an asymptotic quantum number—precisely the Dubna narrative.
@@
This structuralist stance treats physical theories not as descriptions of fundamental entities but as networks of restrictions that generate observable phenomena—a view that finds natural expression in both the bootstrap program and the Dubna approach to color.
@@
The USSR and pre-USSR philosophical tradition includes holist frameworks that, while not direct influences, conceptually resonate with structural approaches in physics. These frameworks emphasize explanation via global restrictions, symmetry, and system-level organization—features aligned with the Dubna school’s early non-dynamical conception of quark color. For example, Vladimir Vernadsky treated the biosphere as a planetary-scale system whose organizing principles govern local dynamics. His notion of the noosphere—human thought as a geophysical force—anticipates modern views of scale-dependent structure. Vernadsky’s framework aligns with effective field theory and renormalization group approaches in its holist orientation: macro-level patterns exhibit autonomy from micro-details, and explanation centers on the global constraint architecture that renders large-scale behavior intelligible [58].
These philosophical traditions should not be read as methodological prescriptions that directly determined Dubna’s scientific approach. An example is the Bogolyubov–Shirkov renormalization group [71], grounded in axiomatic quantum field theory, which reflects a reductionist orientation: it formalizes how local interaction structures discipline behavior across energy scales, aiming to preserve consistency from micro-level dynamics upward—yet its emphasis on symmetry and analyticity also lends it a structural, quasi-holist dimension.
Rather, holist frameworks represent the broader intellectual ecosystem in which Soviet theoretical physics developed during the Cold war. Scientific communities, while governed by empirical and mathematical conditions, nevertheless operate within cultural contexts that can shape methodological preferences and theoretical sensibilities. The presence of these holist traditions in USSR thought may have helped cultivate an intellectual environment where limit-based, structural explanations felt particularly natural—not as a matter of philosophical determinism, but as part of the conceptual resources available to physicists thinking through fundamental problems. This cultural context helps explain why the Dubna approach to color, with its emphasis on structural restrictions over dynamical charges, emerged as a coherent alternative to the gauge-theoretic paradigm developing elsewhere.
@@ 4.1 Problem choice and ontology
The E–36 experiment and its chain were designed to interrogate the high–energy, small–$|t|$ corner of elastic $pp$ scattering where the native theoretical grammar is the Regge/S-matrix one, not perturbative QCD. Already in the early 1970s—and still today for soft elastic diffraction—perturbative QCD offers no general, first-principles description of forward amplitudes, whereas the Regge framework provides a compact language of analytic regularities, pole/trajectory structure, and factorization tailored to this regime. [^1] On the domain-limited tractability of perturbative QCD for small-angle scattering and the Regge/S-matrix idiom used by the E–36 collaboration, see the historical analysis in Phys. in Perspective 18 (2016) 357–378, esp. the section “Experiment Ahead of QCD.” In Regge theory unknown amplitudes are parametrized by Regge trajectories (Reggeons) and their residues; the vacuum exchange that organizes high–energy diffraction is the pomeron.[59, 17, 60, 61, 62, 63] It was precisely this observables-first ontology—analytic $S$ with poles/cuts, crossing, unitarity—that E–36 was built to test.
@@ 4.2 The E36 experimental design as a realization of the S-matrix program
The technological development itself was theory-driven: K.D. Tolstov’s cryogenic target design at JINR emerged from research priorities set by Pomeranchuk’s work on high-energy scattering theory. The target’s optimization for studying elementary proton interactions rather than searching for new particles reflected the S-matrix tradition’s focus on fundamental scattering processes over particle discovery.
@@
While later QCD reinterprets the Pomeron as a color-singlet gluon ladder, the E–36 chain neither requires nor discriminates that ontology; its evidential force lies in validating amplitude-level conditions that the Dubna program treated as primary.
@@ 4.3 Results as S-matrix constraints
E–36’s initial Physical Review Letters reported (a) a rise of $B(s)$ with energy across 88400400 GeV and (b) a transition to positive $\rho$ in the few-hundred GeV range. [^5] PRL slope and $\rho$ papers: Phys. Rev. Lett. 31 (1973) 1088–1091; 31 (1973) 1367–1370. The $\rho>0$ finding around $\sim\!300$ GeV catalyzed a re-reading of the Pomeranchuk theorem among practitioners and clarified how forward analyticity requirements should be applied in practice; [^6] For discussion of the $\rho$ sign change and its interpretive consequences, see [65, 17]. together with the measured shrinkage, these data were absorbed naturally in the Regge phenomenology that motivated the experiment. [^7] On the centrality of Regge/Pomeron language in E–36 analysis and the enduring absence of a first-principles QCD derivation of Reggeons, see [17, 62, 63].
Crucially, the collaboration could—and did—interpret its forward data entirely within the S-matrix/Regge idiom, without invoking quarks or gluons.888“The participants in E–36 and its subsequent chain of experiments did not have to refer to QCD or the quark model in their papers in order to interpret their results.” [64, 65, 66, 67] This was not a deficiency but a principled alignment of domain and ontology: when the task is to delimit $f(s,t)$ near $t=0$, the right language is the one that treats $f$ as primary.
@@
E–36 inhabited the soft, small–$|t|$ domain where QCD’s perturbative “crispness” is inapplicable and where QCD—then nascent—offered neither mechanism nor method for diffraction. In that domain, explanation comes from global structure: the pole content and analyticity of $S$, Regge shrinkage, and the high–energy behavior codified by the pomeron. [^9] On the historical timing (E–36 underway prior to asymptotic freedom) and on the acknowledged gap between soft Regge behavior and QCD derivations, see [67, 66, 10, 11] and [17, 62, 63]. Treating forward conditions as first-class physics—and building an apparatus around them—is non-reductionist in the precise sense relevant to the S-matrix tradition.